OpenStudy (anonymous):

-12x - 24 A. -6(2x - 6) B. -12(x + 2) C. -12(x - 2) D. -6(2x + 6)

jigglypuff314 (jigglypuff314):

-12x - 24 12 and 24 have factor of 12 in common when you factor out a -12 -12/-12 = 1 -24/-12 = 2 so you'll be left with -12(1x + 2)

OpenStudy (anonymous):

thank you

jigglypuff314 (jigglypuff314):

glad I could help :)